OpenStudy (anonymous):

\[L^{-1} (\frac{2s^3}{s^2-81})\] The actual problem in the problem set is \[L^{-1} (\frac{2s^3}{s^4-81})\]

hartnn (hartnn):

i like Laplace

OpenStudy (anonymous):

\[L^{-1} (\frac{2s^3}{s^2-81})=L^{-1} (2s+\frac{81}{s-9}+\frac{81}{s+9})\]

OpenStudy (anonymous):

I don't like Laplace /_\

hartnn (hartnn):

i assume, thats correct, so whats the problem ?

OpenStudy (anonymous):

\[L^{-1}(2s)=trouble\] Btw.. I... suddenly... discovered that... I read the problem wrong :\

OpenStudy (anonymous):

But how to find \(L^{-1}(2s)\)?

OpenStudy (anonymous):

For the correct question in the problem set: \[L^{-1} (\frac{2s^3}{s^4-81})\]\[=L^{-1} (\frac{2s^3}{s^4-81})\]\[=L^{-1} (\frac{s}{s^2+9}+\frac{1}{2(s-3)}+\frac{1}{2(s+3)})\]\[=cos(3t)+cosh(3t)\]

hartnn (hartnn):

sorry, my computer restarted.... to get L^{-1} (2s), u just replace 's' by 2s ....

OpenStudy (anonymous):

L^{-1} (s) also looks weird to me... Usually, it's a/s^n which I can do the inverse easily (easier, I mean)..

hartnn (hartnn):

ohh...i read that wrong .. u know whats L^{-1} s ?

hartnn (hartnn):

its \(\delta'(t)\) so, \(L^{-1}(2s) = 2\delta'(t)\)

OpenStudy (anonymous):

What is \(\delta'(t)\)?

hartnn (hartnn):

u know what is \(\delta(t) ?\)

hartnn (hartnn):

Dirac delta function.. and ' means its derivative

OpenStudy (anonymous):

I.. don't remember I've learnt such function...

hartnn (hartnn):

really ? http://en.wikipedia.org/wiki/Dirac_delta_function 0 everywhere except origin.....like an impulse.

OpenStudy (anonymous):

I think I've only learnt int. from 0 to infty and from s to infty case, not the -ve infty to +ve infty... Thanks for introducing a new friend to me though :)

hartnn (hartnn):

hmm... welcome ^_^

OpenStudy (anonymous):

I want to get back my words. I don't know the name of this new friend, but he's just an old friend of mine :'( I'm sorry!!